# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

### v1

@@ Abstract
While recent multilingual automatic speech recognition models claim to support thousands of languages, ASR for low-resource languages remains highly unreliable due to limited bimodal speech and text training data. Better multilingual spoken language understanding (SLU) can strengthen massively the robustness of multilingual ASR by levering language semantics to compensate for scarce training data, such as disambiguating utterances via context or exploiting semantic similarities across languages. Even more so, SLU is indispensable for inclusive speech technology in roughly half of all living languages that lack a formal writing system. However, the evaluation of multilingual SLU remains limited to shallower tasks such as intent classification or language identification. To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses topical speech classification in 102 languages and multiple-choice question answering through listening comprehension in 92 languages. We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU. Our results show that cascaded systems exhibit greater robustness in multilingual SLU tasks, though speech encoders can achieve competitive performance in topical speech classification when appropriately pre-trained. We further find a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, highlighting the mutual benefits between acoustic and semantic speech representations.111The datasets are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
@@
Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila - Quebec AI Institute {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
@@ 1 Introduction
State-of-the-art multilingual automatic speech recognition (ASR) systems claim to support thousands of languages Pratap et al. (2024); Chen et al. (2024). However, these systems still struggle to reliably transcribe low-resource languages due to the scarcity of bimodal speech and text data in those languages.
Massively multilingual ASR models are commonly pre-trained on the self-supervised wav2vec-BERT 2.0 objective Baevski et al. (2020). Such pre-training learns acoustic features that align representations with phonetics rather than semantics of spoken language Choi et al. (2024).
We therefore argue that speech models require strong multilingual spoken language understanding (SLU) capabilities to enhance massively multilingual ASR for low-resource languages.
Multilingual speech models better at semantic encoding of speech tend to be more robust in ASR too, as they utilize context-related cues, leverage cross-lingual similarities (e.g., cognates, shared subwords, or similar syntax), and better support code-switching in mixed-language speech (cf. §5). Therefore, multilingual SLU capabilities are a key feature of speech models that enable sample-efficient cross-lingual transfer to low-resource or entirely unsupported languages.
Moreover, SLU is a crucial component of inclusive speech technology for languages without a formal writing system, as these languages cannot rely on coupling ASR with language models to offload semantic understanding. This limitation affects approximately half of all living languages.
@@
Despite all this, existing multilingual datasets for evaluating the SLU of speech models have significant limitations. The Minds14 benchmark assesses SLU on 14 high-resource languages for intent classification, which often only requires identifying shallow semantics, such as detecting specific keywords Gerz et al. (2021). SpeechTaxi comprises spoken topical classification for Bible verses in 28 diverse languages, and as such does poorly corresponds to real-world usage scenarios Keller and Glavaš (2024).
To address the shortcomings of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores). We first filter out silent and noisy instances from Fleurs and then map the remaining data back to Flores Team et al. (2022).
@@
The resulting dataset is then merged with SIB-200 Adelani et al. (2024) and Belebele Bandarkar et al. (2023) to create a topical utterance classification benchmark for 102 languages and a multiple-choice question answering on spoken paragraphs for 92 languages, respectively.
@@
Contributions. 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mse). We extensively benchmark state-of-the-art speech models on multilingual semantic speech classification and multiple-choice from spoken paragraphs. Like prior work, we observe that Cascaded Systems (cs) remain the more robust option Keller and Glavaš (2024). We are the first to show that mse pre-trained on language understanding objectives can yield performance competitive to that of cs on topical speech classification. We further isolate utterance quality as a factor in SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can non-negligibly deteriorate transfer performance. 2) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT). This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust. 3) We ablate speech classification on SeamlessM4Tv2-Large Seamless Communication et al. (2023), comparing the performance of the model with and without the final length adaptor layer. We observe that performance decreases only slightly when accounting for the reduction in model size. This ablation further underscores that pre-training objectives, rather than architectural choices, drive multilingual SLU and, through SLU, the overall capabilities of multilingual speech technology.
@@ 2 Related Work
Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training. Among these models, mHubert employs the self-supervised wav2vec-BERT 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features Boito et al. (2024). MMS-1b is also pretrained with wav2vec-BERT 2.0 on an large corpus of 491K hours of speech spanning 1,406 languages Pratap et al. (2024). Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and VAD on 680k hours of audio Radford et al. (2022). SeamlessM4Tv2-Large is a multilingual multimodal translation model Seamless Communication et al. (2023). It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with wav2vec-BERT 2.0 on 4.5M hours of audio. The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives. The authors perform KD from the text encoder to the speech encoder by minimizing the KL-divergence between the decoder’s token output distributions on bi-modal speech and text data.
Multilingual SLU. The evaluation of multilingual Speech Language Understanding (SLU) is constrained to a limited set of tasks. SLU has been significantly shaped by task-oriented dialogue (ToD), leading to the creation of datasets focused on ToD-specific tasks such as intent classification and slot filling. These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords. Additionally, commonly used utterance-level SLU tasks like language identification and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encoding. As a result, the majority of SLU datasets are predominantly in English. Multilingual exceptions are otherwise limited. The Minds14 dataset for intent classification only includes 14 high-resource languages Gerz et al. (2021). SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains Keller and Glavaš (2024).
Concurrent Work. Costa-jussà et al. (2024) concurrently released the multilingual speech benchmark 2M-Belebele for 74 languages. In this dataset, the authors first extend Flores by ca. 20% by incorporating human recordings for both sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele. Costa-jussà et al. (2024) find that a cs-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by ca. 8 percentage points (pp).
@@ 3 Fleurs-SLU
In Fleurs-SLU, we postprocess and joins datasets that are based on Flores to create a massively multilingual SLU benchmark for speech classification and multiple-choice question answering from spoken paragraphs.
@@ 3.1 Core Datasets
Flores. Flores is a machine translation benchmark consisting of 3,001 sentences extracted from English Wikipedia paragraphs, which have been professionally translated into over 200 languages Team et al. (2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets. The authors did not release the test set.
Fleurs. Fleurs comprises 2.3 spoken utterances, on average, per sentence from the public dev and devtest splits of 102 languages in Flores Conneau et al. (2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores. Fleurs is used to evaluate multilingual ASR, language identification, as well as speech-to-text and text-to-speech translation in all language directions.
@@
Belebele. Belebele is a multiple-choice machine reading comprehension benchmark parallel across 122 language variants Bandarkar et al. (2024). The authors first reconstruct paragraphs from Flores sentences. Four multiple-choice questions are created per English paragraph and then professionally translated into 121 language variants. Belebele comprises 900 questions, each associated with one of 488 distinct passages.
@@ 3.2 Benchmark Creation
1) Merging Fleurs & Flores. We begin by removing silent instances from Fleurs.555The numbers of removed examples are listed in Appendix A.6. We first normalize the loudness to an RMS target level of . We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if voice activity is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect. Lastly, we manually verify our approach on 50 randomly sampled predicted silent instances. We find only one borderline misclassified example, which is very noisy but still comprehensible. We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
2a) SIB-Fleurs. For each language, we pool instances from the training, validation, and test splits of SIB-200 Adelani et al. (2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before. The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset. This segmentation ensures compatibility for speech models that may be trained on ASR using the training set of Fleurs prior to SIB-Fleurs evaluation. This also ensures that the speakers in the training set are different from those in the validation and test sets. Our final dataset comprises a median of 728 instances for training, 70 for validation, and 174 for testing across 102 languages. Figure 1 displays the available samples by split per language.999The full data is available in Appendix A.2.
2b) Belebele-Fleurs. We merge our Fleurs-Flores sentences with Belebele paragraphs by intersecting the URLs of the texts. We discard all paragraphs that are not complete in Fleurs-Flores. We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters). Belebele-Fleurs contains, on average, 709 evaluation instances (with a median of 772) across 92 language variants. Figure 2 illustrates the number of samples recovered from the original 900 paragraphs by language variant.
@@ 4.1 Tasks and Languages
Belebele-Fleurs. For cs, we fine-tune and validate models on the English training and development sets compiled by Bandarkar et al. (2024), respectively. We jointly embed the paragraph, question, and choices with text encoders. We then average the token representations of each choice and project the resulting embedding via head to a logit . We minimize the cross-entropy of the concatenated choice logits to the label choice.101010We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input. We leave such evaluation to future work.
@@ 4.2 Cross-Lingual Transfer Setups
Speech Classification. We evaluate state-of-the-art mse for speech classification that differ in pre-training objectives and training data (cf. §2). We include MMS-1b without intermediate fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on a broad range of multilingual datasets (‘MMS-1b-all’). For ASR fine-tuning, Pratap et al. (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen. For task fine-tuning, we freeze the language adapters to facilitate zs-xlt. We further evaluate on the speech encoders of Whisper-v3-Large and SeamlessM4Tv2-Large.
Cascading. Cascaded systems (cs) perform xlt in two steps. cs first transcribe speech into text with an ASR model. The transcription is then fed into an text encoder paired with a classification head. We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large. Note that speech-to-text-translation into English (S2ETT) with cs is equivalent to TTest as we merely monolingual LLMs are needed for xlt. We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024). For SIB-Fleurs, we base our cs experiments on Roberta-Large Liu et al. (2019).111111We did not find improvements by using LLMs in SIB-Fleurs and therefore run broad-spectrum experiments with the more efficient Roberta-Large. For Belebele-Fleurs, we run our experiments on LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024). LLM2Vec BehnamGhader et al. (2024) enables bidirectional attention and continues training on both masked next-token prediction and SimCSE Gao et al. (2021). NLLB-LLM2Vec fuses LLM2Vec with the encoder of the multilingual NLLB translation model to extend its NLU capabilities to more than 200 languages Schmidt et al. (2024).

### v2

@@ Abstract
Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system, since these languages cannot pair automatic speech recognition (ASR) with language models to benefit from language technology. Even if low-resource languages possess a writing system, ASR for these languages remains unreliable due to limited bimodal speech and text training data. Better SLU can strengthen the robustness of massively multilingual ASR by levering language semantics to disambiguate utterances via context or exploiting semantic similarities across languages. However, the evaluation of multilingual SLU remains limited to shallow tasks such as intent classification or language identification. To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering through listening comprehension spanning 944 hours of speech across 92 languages. We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU. Our results show that cascaded systems exhibit greater robustness in multilingual SLU tasks, though speech encoders can achieve competitive performance in topical speech classification when appropriately pre-trained. We further find a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, highlighting the mutual benefits between acoustic and semantic speech representations.111The datasets are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
@@
Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila, McGill University and Canada CIFAR AI Chair {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
@@ 1 Introduction
Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology Ethnologue (2017). Spoken language understanding (SLU) is hence a critical feature of inclusive technology for these languages, which cannot rely on combining automatic speech recognition (ASR) with language models to offload semantic speech understanding. However, the datasets currently available for evaluating SLU in truly relevant languages suffer from significant limitations. For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords Gerz et al. (2021). Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage Keller and Glavaš (2024).
Moreover, even when a low-resource language has a writing system, multilingual ASR often struggles to reliably transcribe that language due to the limited availability of bimodal speech and text data. Pre-trained on the self-supervised wav2vec-BERT objective Baevski et al. (2020), these models embed acoustic features that capture phonetic rather than semantic information Choi et al. (2024). Conversely, multilingual speech models that excel in semantic encoding tend to be more robust in ASR tasks by, e.g., leveraging context-related cues or cross-lingual similarities such as cognates or similar syntactic structures (cf. §5.1). This underscores that robust SLU is an instrumental component of genuinely inclusive speech technology.
@@
To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores). We first filter out silent instances from Fleurs Conneau et al. (2022) and then map the remaining data back to Flores Team et al. (2022). The resulting dataset is then merged with SIB-200 Adelani et al. (2024) and Belebele Bandarkar et al. (2023) to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
@@
Contributions. 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mSE). SIB-Fleurs supports semantic utterance classification with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks. Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages Costa-jussà et al. (2024). 2) We extensively benchmark state-of-the-art speech models on SIB-Fleurs and Belebele-Fleurs. Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding Keller and Glavaš (2024). We are the first to show that mSE pre-trained on language understanding objectives can yield performance competitive to that of CS on speech classification. We further isolate utterance quality as an important factor in multilingual SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can substantially deteriorate transfer performance. 3) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT). This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust.
@@ 2 Related Work
Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training. Among these models, mHubert employs the self-supervised wav2vec-BERT 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features Boito et al. (2024). MMS-1b is also pretrained with wav2vec-BERT 2.0 on an large corpus of 491K hours of speech spanning 1,406 languages Pratap et al. (2024). Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio Radford et al. (2022). SeamlessM4Tv2-Large is a multilingual multimodal translation model Seamless Communication et al. (2023). It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with wav2vec-BERT 2.0 on 4.5M hours of audio. The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives. The authors perform KD from the text encoder to the speech encoder by minimizing the KL-divergence between the decoder’s token output distributions on bi-modal speech and text data.
Multilingual SLU. The evaluation of multilingual SLU is constrained to a limited set of tasks. SLU has been significantly shaped by task-oriented dialogue (ToD), with datasets created focusing on ToD-specific tasks such as intent classification and slot filling. These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords. Additionally, commonly used utterance-level SLU tasks like language identification (LID) and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encodings. As a result, the majority of SLU datasets are predominantly in English. Multilingual exceptions are otherwise limited. The Minds14 dataset for intent classification only includes 14 high-resource languages Gerz et al. (2021). SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains Keller and Glavaš (2024).
Concurrent Work. Costa-jussà et al. (2024) concurrently released the multilingual SLU benchmark 2M-Belebele for 74 languages. In this dataset, the authors first extend Flores by ca. 20% by incorporating human recordings for both sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele. Costa-jussà et al. (2024) find that a CS-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by ca. 8 percentage points (pp).
@@ 3 Fleurs-SLU
We postprocess datasets that are based on Flores to create ‘Fleurs-SLU’, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs.
@@ 3.1 Core Datasets
Flores. The authors professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation Team et al. (2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets. The authors did not release the test set.
Fleurs. Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores Conneau et al. (2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores. Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech translation in all language directions.
@@
Belebele. Belebele is a multiple-choice reading comprehension benchmark for 122 languages Bandarkar et al. (2024). The authors reconstruct paragraphs from Flores sentences. They generate 1–2 questions per English paragraph, which are professionally translated into 121 languages. Belebele comprises 900 questions spanning 488 passages.
@@ 3.2 Benchmark Creation
1) Merging Fleurs & Flores. We begin by removing silent instances from Fleurs.555Counts of removed examples are listed in Appendix A.6. We first normalize the loudness to an RMS target level of . We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect. Lastly, we verify our approach on 50 randomly sampled predicted silent instances. We find only one borderline misclassified example, which is very noisy but still comprehensible. We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
2a) SIB-Fleurs. For each language, we pool instances from the training, validation, and test splits of SIB-200 Adelani et al. (2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before. The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset. This segmentation ensures compatibility for speech models that may be trained on ASR using the training set of Fleurs prior to SIB-Fleurs evaluation. This also ensures that the speakers in the training set are different from those in the validation and test sets. Table 2 lists aggregated statistics on the instance- and the utterance-level for SIB-Fleurs. Appendix A.2 provides a list of samples by split per language.
| Task | Utterance |
| Classification | |
| Languages | 102 |
| Classes (Topics) | 7 |
| Utterances per sample | 2.2 (2.0) |
| Duration per utterance (s) | 12.9 (11.6) |
| Total audio (hr) | 692 |
| Samples by Split | |
| Training | 696 (728) |
| Validation | 66 (70) |
| Test | 163 (174) |
2b) Belebele-Fleurs. We merge our Fleurs-Flores sentences with Belebele paragraphs by intersecting the URLs of the texts. We discard all paragraphs that are not complete in Fleurs-Flores. We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters). Table 2 provides a detailed summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
Task Multiple-Choice QA Spoken Paragraphs Languages 92 Answer Choices 4 Questions per paragraph 1.8 (1.8) Sentences per paragraph 3.6 (4.0) Utterances per sentence 2.0 (2.0) Duration per utterance (s) 12.6 (12.0) Duration per paragraph (s) 47.0 (43.4) Total audio (hr) 944 Samples (Paragraphs) 709 (771)
@@ 4.1 Tasks and Languages
Belebele-Fleurs. We train and validate models on the English training and dev sets of Bandarkar et al. (2024), respectively. We jointly embed the paragraph, question, and choices with text encoders. We then average the token encodings of each choice and project the choice embedding via head to a logit . We minimize the cross-entropy of the concatenated choice logits to the label choice.999We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input. We leave such evaluation to future work.
@@ 4.2 Cross-Lingual Transfer Setups
Speech Classification. We evaluate state-of-the-art mSE for speech classification (cf. §2). We include MMS-1b without fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1b-all’). For ASR fine-tuning, Pratap et al. (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen. For task fine-tuning, we freeze the adapters to facilitate zs-xlt. We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large.
Cascading & Text. Cascaded systems (CS) perform xlt in two steps. CS first transcribe speech into text with an ASR model. The transcription is then fed into an text encoder paired with a classification head. We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large. Note that speech-to-text-translation into English (S2ETT) with CS is equivalent to TTest as merely monolingual LLMs are needed for xlt. We further evaluate on the original text of SIB and Bebelebe, which we translate to English with SeamlessM4Tv2-Large.101010We use the gold Q&A for the respective language (i.e., English for ). Otherwise, we would need to interleave MT models, e.g., S2ETT of paragraphs with Whisper, but translating textual Q&A with SeamlessM4Tv2. We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024). We run our text-based and CS experiments on Roberta-Large Liu et al. (2019) as well as LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024). LLM2Vec BehnamGhader et al. (2024) enables bidirectional attention and continues training on both masked next-token prediction and SimCSE Gao et al. (2021). NLLB-LLM2Vec fuses LLM2Vec with the encoder of the NLLB translation model to extend its NLU capabilities to 200 languages Schmidt et al. (2024).